OpenStudy (flvskidd):

In your lab, a substance's temperature has been observed to follow the function T(x) = (x − 4)3 + 6. The turning point of the graph is where the substance changes from a liquid to a gas. Using complete sentences in your written answer, explain to your fellow scientists how to find the turning point of this function. Hint: The turning point of the graph is similar to the vertex of a quadratic function.

OpenStudy (alekos):

That would be where the derivative equals zero

OpenStudy (flvskidd):

is that my answer?

OpenStudy (anonymous):

Okay. I think the a,b, and c are from a x^3 + b x^2 +c x + d unfortunately, you have (x - 4)^3 + 6 which is (x-4) (x-4) (x-4) + 6 we have to multiply that out to get it in the form so we can find a, b and c

OpenStudy (anonymous):

first do (x-4)(x-4)

OpenStudy (anonymous):

Which is \[x^2 - 8x + 16\]

OpenStudy (anonymous):

Now do: \[(x^2 -8x+16)(x-4)\]

OpenStudy (anonymous):

Which is: \[x^3 - 12x^2 + 48x - 64\]

OpenStudy (anonymous):

So your function is: \[(x−4)^3+6=x^3−12x^2+48x−64+6=x^3−12x^2+48x−58\]

OpenStudy (anonymous):

so A=1 B=-12 C=48

OpenStudy (anonymous):

So it the answer is x=12/3 = 4 that is the turning point of this curve Here is a graph:

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @FLVSKidd

OpenStudy (flvskidd):

Thank you Tom. Yes it did help.

OpenStudy (anonymous):

Haha. No problem! @FLVSKidd

OpenStudy (alekos):

that's one way of doing it but........... taking the derivative of the function is much easier. then make it equal to zero to get the turning point